A coffee distributor needs to mix a(n) Kenya coffee blend that normally sells for $8.40 per pound with a Gazebo coffee blend that normally sells for $9.30 per pound to create 50 pounds of a coffee that can sell for $8.98 per pound. How many pounds of each kind of coffee should they mix
Answer by mananth(16946)

Price ---------------- quantity
Kenya coffee 8.40% ---------------- x pounds
Gazebo coffee 9.30% ------ 50 - x pounds
Mixture 8.98% ---------------- 50
Total 50 pounds
8.40% x + 9.30% ( 50 - x ) = 8.98% * 50
8.4 x + 9.3 ( 50 - x ) = 449
8.4 x + 465 - 9.3 x = 449
8.4 x - 9.3 x = 449 - -465
-0.9 x = -16
/ -0.9
x = 17.78 pounds 8.40% Kenya coffee
32.22 pounds 9.30% Gazebo coffee

m.ananth@hotmail.ca